DIRECTIONS: USE THE STEPS FOR SOLVING A SYSTEM OF LINEAR EQUATIONS OR INEQUALITIES TO ANSWER THE QUESTIONS.....

1.) Sam paddles his canoe upstream, AGAINST the current, for 12 miles. It takes him 2 hours. Coming back downstream, WITH the current, he covers the same distance in only an hour. How fast can Sam paddle his canoe in still water? How fast is the current?

Answer by stanbon(48558)

Sam paddles his canoe upstream, AGAINST the current, for 12 miles. It takes him 2 hours. Coming back downstream, WITH the current, he covers the same distance in only an hour. How fast can Sam paddle his canoe in still water? How fast is the current?
--------------
Let the current speed be "c"
Let the boat speed be "b"
Speed = distance/time
------------
EQUATIONS:
b-c=12/2
b+c=12/1
---------
b-c=6
b+c=12
----------
Add the two equations to get:
2b=18
b=9 mph (speed of the boat in still water)
Then, b+c=12; c=3 mph (speed of the current)
================
Cheers,
Stan H.